Solve the following system of equations.
x - 2y = 5
2x - 4y = 10

Answer by ewatrrr(24785)

 
Hi,
Solve the following system of equations.
x - 2y = 5
2x - 4y = 10
eliminating y by multiplying 1st EQ thru by -2
-2x + 4y = -10
2x - 4y = 10
0 = 0 There are an infinitely amount of solutions to thid system
Graphically they are the same line.